How do you the sum of the first 850 odd numbers.
Answer by stanbon(75887)

How do you the sum of the first 850 odd numbers?
------------
The 1st odd digit is one: a(1) = 1
The difference between each odd digit is d = 2
----
To find the nth term of the sequence use a(n) = a(1) + (n-1)d
a(850) = 1+849(2) = 1699
------------------------------
The sum of n terms of n terms of the sequance is s(n) = (n/2)(a(1)+a(n))
S(850) = (850/2)(1+1699+1)
S(850) = 425*850 = 361,250
=============================
Cheers,
Stan H.
=============